# MENTOR: Efficient Multimodal-Conditioned Tuning for Autoregressive Vision Generation Models

### v1

@@ 5 Conclusion
In this work, we introduced a controllable and efficient autoregressive framework for complex multimodal image generation, offering a compelling alternative to diffusion-based methods. By unifying multimodal inputs within an AR model and leveraging a two-stage training paradigm, our method achieves state-of-the-art performance on challenging benchmarks—despite a modest model size, suboptimal base component, and limited training resources. These results underscore efficiency, scalability, and controllability of our method, establishing it as a efficient foundation for building versatile, fine-grained visual generation systems capable of handling complex multimodal prompts.
Appendix A Technical Appendices and Supplementary Material
This Supplementary Material is organized as follows.
@@ Appendix A Technical Appendices and Supplementary Material
  • In Appendix B, we provide the training details of Mentor, including initialization (§ B.1), training procedures (§ B.2), and multi-image training strategy(§ B.3).
@@
  • In Appendix C, we show quantitative evaluations of our method on text-to-image generation benchmarks.
@@
  • In Appendix D, we detail our data construction pipeline and the dataset details used across the two-stage training.
@@
  • In Appendix E, we elaborate on the experimental setup, including datasets and metrics (§ E.1), as well as detailed descriptions of baseline methods(§ E.2).
@@
  • In Appendix F, we present qualitative results that demonstrate the capabilities of Mentor in various settings, such as image reconstruction(§ F.1), segmentation(§ F.2), multi-image generation(§ F.3), and in-context image generation(§ F.4).
@@
  • In Appendix G, we demonstrate the versatility of Mentor across diverse multimodal generation tasks, including segmentation, subject-driven generation, and multimodal in-context learning.
  • In Appendix H, we discuss the current limitations of our method, such as its dependence on autoregressive generators, generation fidelity, and safety considerations.
@@ B.1 Initialization Details
The multimodal encoder is initialized using the vision encoder from CLIP-Large-Patch14 [Radford et al., 2021], with an image receptive field of $224\times 224$, and the FlanT5-XL encoder [Chung et al., 2022], with a context length of 512 tokens. This encoder converts each image into 256 tokens for use as context in the generator.
To implement the MLP-based projection, we train the MLP projector on the LLaVA-CC3M-Pretrain-595K dataset [Liu et al., 2023], following the alignment training setup used by LLaVA. Specifically, we freeze both the vision and text encoders (CLIP-Large-Patch14 and FlanT5-XL, respectively) and train only the MLP layers. The resulting pretrained MLP layers are then directly incorporated into the multimodal encoder of Mentor.
@@
The projector consists of a two-layer MLP with an intermediate dimension of 4,096, employing SiLU activation functions. The autoregressive generator is initialized from LlamaGen-XL [Sun et al., 2024b] with 775 million parameters. However, the original LlamaGen implementation contains a fundamental error in its 2D Rotary Positional Embedding[Lu et al., 2023, Fang et al., 2024] (ROPE) mechanism***https://github.com/FoundationVision/LlamaGen/issues/54, which leads to a loss of information in the query and key vectors during attention computation. To address this, we correct the ROPE implementation in our code and continue training the revised model on both the Midjourney dataset [Emporium, 2024] and the LAION-COCO dataset used in LlamaGen pretraining, effectively replicating the original pretraining conditions. This continued training enables the model to adapt to the corrected ROPE mechanism. The resulting model is then used to initialize our autoregressive generator.
@@ B.3 Multi-Image Training
In multi-image training scenario, the context length of Mentor is expanded to 1,280 tokens to accommodate up to 4 images per context. For the Query-based variant of Mentor, token compression techniques enable processing up to 14 images per context with 512 context length.
@@
Qualitative assessments, presented in Figure 6, demonstrate that multi-image training significantly enhances the model’s capability to preserve detailed visual information in complex multimodal contexts.
@@ Appendix C Text-to-Image Generation Evaluation
Autoregressive Chameleon Team [2024a] - - - - - - $0.39$
LWM Liu et al. [2024a] $0.93$ $0.41$ $0.46$ $0.79$ $0.09$ $0.15$ $0.47$
LlamaGen Sun et al. [2024a] $0.71$ $0.34$ $0.21$ $0.58$ $0.07$ $0.04$ $0.32$
Show-o Xie et al. [2024] $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
Emu3333-Gen Wang et al. [2024c] $0.98$ $0.71$ $0.34$ $0.81$ $0.17$ $0.21$ $0.54$
Janus Wu et al. [2024b] $0.97$ $0.68$ $0.30$ $0.84$ $0.46$ $0.42$ $0.61$
Mentor $0.87$ $0.38$ $0.18$ $0.67$ $0.08$ $0.13$ $0.38$
Diffusion LDM Rombach et al. [2022a] $0.92$ $0.29$ $0.23$ $0.70$ $0.02$ $0.05$ $0.37$
SDv1.51.51.51.5 Rombach et al. [2022a] $0.97$ $0.38$ $0.35$ $0.76$ $0.04$ $0.06$ $0.43$
PixArt-$\alpha$ Chen et al. [2023a] $0.98$ $0.50$ $0.44$ $0.80$ $0.08$ $0.07$ $0.48$
SDv2.12.12.12.1 Rombach et al. [2022a] $0.98$ $0.51$ $0.44$ $0.85$ $0.07$ $0.17$ $0.50$
DALL-E 2 Ramesh et al. [2022] $0.94$ $0.66$ $0.49$ $0.77$ $0.10$ $0.19$ $0.52$
SDXL Podell et al. [2024] $0.98$ $0.74$ $0.39$ $0.85$ $0.15$ $0.23$ $0.55$
DALL-E 3 Betker et al. [2023] $0.96$ $0.87$ $0.47$ $0.83$ $0.43$ $0.45$ $0.67$
SDv3 Medium Esser et al. [2024] $0.98$ $0.74$ $0.63$ $0.67$ $0.34$ $0.36$ $0.62$
Flux.1 Dev [BlackForestLabs, 2024] $0.98$ $0.81$ $0.74$ $0.79$ $0.22$ $0.45$ $0.66$
Dream Engine [Chen et al., 2025] $1.00$ $0.94$ $0.64$ $0.81$ $0.27$ $0.49$ $0.69$
SDv3.5 Large Esser et al. [2024] $0.98$ $0.89$ $0.73$ $0.83$ $0.34$ $0.47$ $0.71$
We evaluate the performance of our model on text-to-image (T2I) generation using the MS-COCO [Lin et al., 2014] and GenEval [Ghosh et al., 2024] benchmarks. Results are reported in Table 7 and Table 6.
@@
Nevertheless, Mentor demonstrates strong performance in multimodal image generation tasks. Thanks to our proposed autoregressive architecture and a two-stage multimodal-conditioned tuning strategy, Mentor effectively integrates both visual and textual modalities during generation. This synergistic design compensates for its weaker generation core, enabling Mentor to surpass more powerful T2I models in multimodal settings, as shown in Table 3. We anticipate that incorporating stronger base generators will further improve performance. Despite its current limitations, our results suggest that Mentor presents a promising and efficient alternative to diffusion-based methods in multimodal scenarios.
Table 7: Zero-shot FID scores on the MS-COCO benchmark. Lower is better.
Model FID $\downarrow$
T2I Models
GLIDE [Nichol et al., 2022] 12.24
Make-A-Scene [Gafni et al., 2022] 11.84
DALL-E 2 [Ramesh et al., 2022] 10.39
SD v1.5 [Rombach et al., 2022c] 9.34
Imagen [Saharia et al., 2022] 7.27
CLIP-Aligned VL2I Models
GILL [Koh et al., 2023] 12.20
Emu [Sun et al., 2023] 11.66
KOSMOS-G [Pan et al., 2024] 10.99
Mentor 19.92
@@ Appendix D Data Construction and Formation
1 Midjourney[Emporium, 2024] Text to Image Generation 700k
Midjourney[Emporium, 2024] Image Reconstruction 180k
@@ 1 Midjourney[Emporium, 2024] Text to Image Generation 700k
2 Midjourney[Emporium, 2024], Synthetic Data Text to Image Generation 600k
@@
Synthetic Data, CC12M[Changpinyo et al., 2021] Image Recovery 150k
Subject200k[Tan et al., 2025] Subject-driven Generation 400k
@@
The dataset is constructed from a combination of open-source resources, such as Midjourney [Emporium, 2024] and CC12M [Changpinyo et al., 2021], along with synthetic data generated via publicly available text-to-image (T2I) models, including Flux.1 [BlackForestLabs, 2024] and Stable Diffusion v3.5 [Esser et al., 2024].
Stage 1 focuses on establishing foundational multimodal alignment capabilities. Specifically, it includes 700k T2I samples from Midjourney [Emporium, 2024], 180k image reconstruction samples also from Midjourney, and 1.6M object segmentation samples generated through our pipeline.
Stage 2 fine-tunes the model with 1.3 million samples. This includes 600k T2I samples—200k from Midjourney and 400k synthesized using open-source T2I models such as Flux.1 [BlackForestLabs, 2024] and Stable Diffusion v3.5 [Esser et al., 2024]. Additionally, we include 150k object segmentation samples and 150k image recovery samples, all derived from synthetic data using segmentation masks. Background images for the image recovery task are randomly selected from CC12M [Changpinyo et al., 2021].
We further incorporate 400k subject-driven image generation samples from Subject200k [Tan et al., 2025]. These samples are re-captioned using Qwen2-VL [Wang et al., 2024a] to extract subject-relevant text and generate comprehensive image descriptions. To enrich the training set, we reverse the input-output image pairs, effectively doubling the usable data to 400k samples.

### v2

@@ 5 Conclusion
In this work, we introduced a controllable and efficient autoregressive framework for complex multimodal image generation, offering a compelling alternative to diffusion-based methods. By unifying multimodal inputs within an AR model and leveraging a two-stage training paradigm, our method achieves state-of-the-art performance on challenging benchmarks—despite a modest model size, suboptimal base component, and limited training resources. These results underscore the efficiency, scalability, and controllability of our method, establishing it as an efficient foundation for building versatile, fine-grained visual generation systems capable of handling complex multimodal prompts.
Limitations
Our work introduces a resource-efficient framework for multimodal-conditioned image generation, primarily aimed at exploring how to balance multimodal guidance for controllable generation rather than maximizing absolute generation quality. Accordingly, our evaluation focuses on verifying the effectiveness of the proposed training paradigm and architectural design under limited-resource settings, using models of comparable capacity and scale. However, the current performance of our approach is inherently constrained by the limitations of generative backbone models. Unlike recent unified or large-scale models trained with thousands of GPUs or on extensive datasets, our experiments are conducted on a smaller scale using publicly available resources. Consequently, Mentor exhibits shortcomings in text-to-image generation, including spatial reasoning, object counting, fine-grained human rendering, and stylization. These limitations reflect the current gap between current SOTA diffusion and autoregressive architectures in terms of generation fidelity and domain generalization. Future work can readily extend this framework by replacing backbone modules and scaling data resources to further enhance image quality. Additionally, while our training data is sourced from publicly available datasets and our synthetic data pipeline includes NSFW safeguards, a comprehensive evaluation of safety, fairness, and potential misuse remains lacking. Future work should incorporate thorough assessments of model biases and unintended behaviors. Finally, while our framework demonstrates strong versatility across diverse multimodal tasks, achieving competitive performance in specific domains may require more specialized training and the integration of more powerful multimodal encoders and generators.
Appendix A Appendix
This Appendix is organized as follows.
  • In §˜A.1, we describe the autoregressive training objective for our framework.
@@ Appendix A Appendix
  • In Appendix˜B, we provide the training details of Mentor, including initialization (§˜B.1), training procedures (§˜B.2), multi-image training strategy (§˜B.3), and RL training with GRPO (§˜B.4).
@@
  • In Appendix˜C, we show quantitative evaluations of our method on text-to-image generation benchmarks.
@@
  • In Appendix˜D, we detail our data construction pipeline and the dataset details used across the two-stage training.
@@
  • In Appendix˜E, we elaborate on the experimental setup, including datasets and metrics (§˜E.1), as well as detailed descriptions of baseline methods (§˜E.2).
@@
  • In Appendix˜F, we present qualitative results that demonstrate the capabilities of Mentor in various settings, such as image reconstruction (§˜F.1), segmentation (§˜F.2), multi-image generation (§˜F.3), and in-context image generation (§˜F.4).
@@
  • In Appendix˜G, we demonstrate the versatility of Mentor across diverse multimodal generation tasks, including segmentation, subject-driven generation, and multimodal in-context learning.
  • In Appendix˜H, we present the human evaluation study validating the reliability of GPT-based metrics.
  • In Appendix˜I, we analyze CP–PF trade-offs via task weight and CFG sweeps.
A.1 Preliminary
Training Objective Our model employs teacher forcing to predict image tokens, conditioned on (i) previously generated tokens and (ii) multimodal context $\mathbf{h}$. Given the multimodal condition: $\mathbf{c}^{(0)}=\{\mathcal{I},\mathcal{T}\}$ (visual and textual inputs), a multimodal encoder $\phi$ first encodes $\mathbf{c}^{(0)}$ and subsequently uses an MLP layer to project them into the space of the image decoder to form a unified representation $\mathbf{h}$:
$\mathbf{H}=\text{MLP}(\phi\bigl(\mathbf{c}^{(0)}\bigr))=(\mathbf{h}_{1},\dots,\mathbf{h}_{M})\in\mathbb{R}^{M\times d},\mathbf{h}_{j}\in\mathbb{R}^{d}.$ (1)
$\theta(\mathbf{y}\mid\mathbf{H})=\prod_{i=1}^{L}\theta\bigl(y_{i}\mid y_{<i},\mathbf{H}\bigr).$ (2)
$\mathcal{L}_{\text{CE}}(\theta,\phi)=-\mathbb{E}_{(\mathbf{y},\mathbf{c}^{(0)})\sim\mathcal{D}}\left[\sum_{i=1}^{L}\log\theta\bigl(y_{i}\mid y_{<i},\mathbf{H}\bigr)\right].$ (3)
Classifier-free Guidance To enhance multimodal generation controllability, we apply Classifier-Free Guidance (CFG) (Sun et al., 2024a). During training, multimodal conditioning $\mathbf{H}$ is replaced by a learned unconditional embedding $\mathbf{H}_{u}$ with probability $p$. At inference time, token logits $\ell_{g}$ are recalculated by interpolating between the conditional logits $\ell_{c}$ (from $\mathbf{H}$) and unconditional logits $\ell_{u}$ (from $\mathbf{H}_{u}$), controlled by a scaling parameter $\lambda$: $\ell_{g}=\ell_{u}+(\ell_{c}-\ell_{u})\times\lambda$.
@@ B.1 Initialization Details
The multimodal encoder is initialized using the vision encoder from CLIP-Large-Patch14 (Radford et al., 2021), with an image receptive field of $224\times 224$, and the FlanT5-XL encoder (Chung et al., 2022), with a context length of 512 tokens. This encoder converts each image into 256 tokens for use as context in the generator.
To implement the MLP-based projection, we train the MLP projector on the LLaVA-CC3M-Pretrain-595K dataset (Liu et al., 2023), following the alignment training setup used by LLaVA. Specifically, we freeze both the vision and text encoders (CLIP-Large-Patch14 and FlanT5-XL, respectively) and train only the MLP layers. The resulting pretrained MLP layers are then directly incorporated into the multimodal encoder of Mentor.
@@
The projector consists of a two-layer MLP with an intermediate dimension of 4,096, employing SiLU activation functions. The autoregressive generator is initialized from LlamaGen-XL (Sun et al., 2024a) with 775 million parameters. However, the original LlamaGen implementation contains a fundamental error in its 2D Rotary Positional Embedding (Lu et al., 2023; Fang et al., 2024) (ROPE) mechanism***https://github.com/FoundationVision/LlamaGen/issues/54, which leads to a loss of information in the query and key vectors during attention computation. To address this, we correct the ROPE implementation in our code and continue training the revised model on both the Midjourney dataset (Emporium, 2024) and the LAION-COCO dataset used in LlamaGen pretraining, effectively replicating the original pretraining conditions. This continued training enables the model to adapt to the corrected ROPE mechanism. The resulting model is then used to initialize our autoregressive generator.
@@ B.2 Training Procedure
Figure 6: Qualitative examples of different methods compared to Mentor on DreamBench++.
@@ B.3 Multi-Image Training
In the multi-image training scenario, the context length of Mentor is expanded to 1,280 tokens to accommodate up to 4 images per context. For the Query-based variant of Mentor, token compression techniques enable processing up to 14 images per context with 512 context length.
@@
Qualitative assessments, presented in Figure˜7, demonstrate that multi-image training significantly enhances the model’s capability to preserve detailed visual information in complex multimodal contexts.
B.4 GRPO Fine-Tuning for Autoregressive Multimodal Generation
To explore reinforcement learning for controllable multimodal generation, we adopt the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) under the autoregressive (AR) training paradigm, following the AR-GRPO setup (Yuan et al., 2025). Given the autoregressive nature of our framework, this enables token-level optimization of generation behaviors, aligning output quality directly with multimodal guidance.
We apply the vanilla GRPO implementation from the trl (von Werra et al., ) library. The model is fine-tuned for 600 steps with a batch size of 12 and a group size of 8, resulting in 96 sampled generations per update. The training corpus includes 73K multimodal samples from UNO-1M (Wu et al., 2025b) and 60K text-to-image samples from BLIP-3O (Chen et al., 2025a). KL regularization is applied toward a frozen reference policy to maintain training stability.
We employ a composite reward that integrates multiple complementary signals reflecting perceptual quality, aesthetic appeal, and semantic fidelity. Specifically, for text-to-image (T2I) generation, the reward combines HPSv2, Aesthetic, PickScore, and QwenVL-7B scores. For multimodal generation, we additionally include a CLIP-image consistency score computed against the reference image. We utilize QwenVL-7B as a unified vision-language reward model to assess generation quality across T2I and multimodal settings.
Our GRPO-based fine-tuning effectively bridges autoregressive modeling and multimodal RL. By integrating structured, interpretable reward signals, we enable direct policy optimization toward higher visual quality and semantic alignment.
(1) Text-to-Image Evaluation Prompt
Instruction: You are a strict reward model for text-to-image (T2I) evaluation. Your goal is to produce an objective visual description and four quantitative sub-scores, with a total score in [0,4]. Inputs: Text prompt: "{PROMPT}" Generated image: <image> Evaluation Procedure: 1. Describe the generated image only — objects, colors, lighting, composition, perspective, and visible defects. Do not infer from the prompt. 2. Provide four sub-scores (each $\in$ [0,1.0]): (A) Prompt/Category Alignment, (B) Completeness & Composition, (C) Realism & Visual Quality, (D) Artifact/Defect Freedom. 3. Compute the final score: $\text{raw}=A+B+C+D$; if any sub-score $\leq$ 0.3, cap final = 1.0. Otherwise, final = raw (rounded to two decimals). Output: JSON with fields {"description": "...", "scores": {"A":..,"B":..,"C":..,"D":..}, "final": ...}
(2) Multimodal Generation Evaluation Prompt
Instruction: You are a strict reward model for subject-driven image generation with a reference image. Use the full [0,5] range and emphasize identity preservation and artifact freedom. Inputs: Reference image: <image> Subject name: "{SUBJECT}" Text prompt: "{PROMPT}" Generated image: <image> Evaluation Procedure: 1. Describe the generated image only, without inference. 2. Provide three sub-scores: (A) Instruction Following (0–0.6), (B) Identity Preservation vs. Reference (0–3.2), (C) Artifact/Defect Freedom (0–1.2). 3. Compute the final score: $\text{raw}=A+B+C$; if A $\leq$ 0.1, B $\leq$ 0.3, or C $\leq$ 0.3, cap final = 1.0. Otherwise, final = raw (rounded to two decimals). Output: JSON with fields {"description": "...", "scores": {"A":..,"B":..,"C":..}, "final": ...}
@@ Appendix C Text-to-Image Generation Evaluation
Autoregressive Chameleon Team (2024b) - - - - - - $0.39$
LWM Liu et al. (2024b) $0.93$ $0.41$ $0.46$ $0.79$ $0.09$ $0.15$ $0.47$
LlamaGen Sun et al. (2024b) $0.71$ $0.34$ $0.21$ $0.58$ $0.07$ $0.04$ $0.32$
Show-o Xie et al. (2024) $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
Emu33-Gen Wang et al. (2024c) $0.98$ $0.71$ $0.34$ $0.81$ $0.17$ $0.21$ $0.54$
Janus Wu et al. (2024b) $0.97$ $0.68$ $0.30$ $0.84$ $0.46$ $0.42$ $0.61$
Mentor $0.87$ $0.47$ $0.16$ $0.65$ $0.11$ $0.17$ $0.40$
Diffusion LDM Rombach et al. (2022a) $0.92$ $0.29$ $0.23$ $0.70$ $0.02$ $0.05$ $0.37$
SDv1.51.5 Rombach et al. (2022a) $0.97$ $0.38$ $0.35$ $0.76$ $0.04$ $0.06$ $0.43$
PixArt-$\alpha$ Chen et al. (2023a) $0.98$ $0.50$ $0.44$ $0.80$ $0.08$ $0.07$ $0.48$
SDv2.12.1 Rombach et al. (2022a) $0.98$ $0.51$ $0.44$ $0.85$ $0.07$ $0.17$ $0.50$
DALL-E 2 Ramesh et al. (2022) $0.94$ $0.66$ $0.49$ $0.77$ $0.10$ $0.19$ $0.52$
SDXL Podell et al. (2024) $0.98$ $0.74$ $0.39$ $0.85$ $0.15$ $0.23$ $0.55$
DALL-E 3 Betker et al. (2023) $0.96$ $0.87$ $0.47$ $0.83$ $0.43$ $0.45$ $0.67$
SDv3 Medium Esser et al. (2024) $0.98$ $0.74$ $0.63$ $0.67$ $0.34$ $0.36$ $0.62$
Flux.1 Dev (BlackForestLabs, 2024) $0.98$ $0.81$ $0.74$ $0.79$ $0.22$ $0.45$ $0.66$
Dream Engine (Chen et al., 2025b) $1.00$ $0.94$ $0.64$ $0.81$ $0.27$ $0.49$ $0.69$
We evaluate the performance of our model on text-to-image (T2I) generation using the GenEval (Ghosh et al., 2024) benchmarks. Results are reported in Table˜6.
@@
Nevertheless, Mentor demonstrates strong performance in multimodal image generation tasks. Thanks to our proposed autoregressive architecture and a two-stage multimodal-conditioned tuning strategy, Mentor effectively integrates both visual and textual modalities during generation. This synergistic design compensates for its weaker generation core, enabling Mentor to surpass more powerful T2I models in multimodal settings, as shown in Table˜3. We anticipate that incorporating stronger base generators will further improve performance. Despite its current limitations, our results suggest that Mentor presents a promising and efficient alternative to diffusion-based methods in multimodal scenarios.
@@ Appendix D Data Construction and Formation
1 Midjourney(Emporium, 2024) Text to Image Generation 700k
Midjourney(Emporium, 2024) Image Reconstruction 180k
@@ 1 Midjourney(Emporium, 2024) Text to Image Generation 700k
2 Midjourney(Emporium, 2024), Synthetic Data Text to Image Generation 600k
@@
Synthetic Data, CC12M(Changpinyo et al., 2021) Image Recovery 150k
Subject200k(Tan et al., 2025) Subject-driven Generation 400k
@@
The dataset is constructed from a combination of open-source resources, such as Midjourney (Emporium, 2024) and CC12M (Changpinyo et al., 2021), along with synthetic data generated via publicly available text-to-image (T2I) models, including Flux.1 (BlackForestLabs, 2024) and Stable Diffusion v3.5 (Esser et al., 2024).
Stage 1 focuses on establishing foundational multimodal alignment capabilities. Specifically, it includes 700k T2I samples from Midjourney (Emporium, 2024), 180k image reconstruction samples also from Midjourney, and 1.6M object segmentation samples generated through our pipeline.
Stage 2 fine-tunes the model with 1.3 million samples. This includes 600k T2I samples—200k from Midjourney and 400k synthesized using open-source T2I models such as Flux.1 (BlackForestLabs, 2024) and Stable Diffusion v3.5 (Esser et al., 2024). Additionally, we include 150k object segmentation samples and 150k image recovery samples, all derived from synthetic data using segmentation masks. Background images for the image recovery task are randomly selected from CC12M (Changpinyo et al., 2021).
We further incorporate 400k subject-driven image generation samples from Subject200k (Tan et al., 2025). These samples are re-captioned using Qwen2-VL (Wang et al., 2024a) to extract subject-relevant text and generate comprehensive image descriptions. To enrich the training set, we reverse the input-output image pairs, effectively doubling the usable data to 400k samples.